# Molecular clouds: do they deserve a non-Gaussian description?

### v1

@@ 2.2 Subsampling and tiling
The presence of physical processes such as gravity in the dynamics of MCs implies spatial variations of their statistical properties, which prevents us from modeling them as stationary stochastic processes. For instance, they usually have strong local overdensities, while being surrounded by diffuse borders. This non-homogeneity makes it difficult to compare such objects as a whole, as well as to properly estimate their statistical properties, when seen as realizations of a random process, as these properties will for instance depend heavily on the identification of the cloud boundary, or on the precise definition of what a cloud is. Furthermore, variance estimates for random processes generally rely on a homogeneity assumption to estimate the intrinsic variability of a process from its spatial variations. More broadly, the estimation of the statistical moments and properties of a process is usually based on spatial averaging, which assumes the homogeneity of the sample.
To avoid this problem, we choose to restrict the comparison to local patches of these MCs, assuming statistical homogeneity within each individual patch. This requires the identification of a characteristic stationarity length over which it can be assumed that the statistical properties of the cloud do not vary significantly. Different patches extracted from the same cloud may have different statistical properties, which will be representative of the non-homogeneity of the cloud as a whole. This encourages us to use small patches. However, an additional difficulty is that these patches must be large enough to make statistical estimates. In particular, variance estimates require sampling beyond the correlation length of the process under study.
In this paper, we have chosen to cut patches with angular sizes ranging from $0.85$° to $2.55$°, that correspond respectively to 333\,3pc and 999\,9pc for a cloud at a typical distance of 200200200\,200pc. We believe this represents a good compromise with regard to the trade-off expressed above. However, we are aware that there may not be an ideal solution, especially as we have no precise estimate of neither the stationarity length nor the correlation length. This means that we cannot rule out some confusion between local intrinsic variability and large-scale variations in the statistical properties of the clouds. However, we believe that, for lack of a better solution, this does not undermine the relevance of this study.
More precisely, we make multiple versions of the original (18 ″ resolution, 3 ″pixel size) column density maps at different pixel sizes (6 ″, 9 ″, 12 ″, and 18 ″). The down-sampling is done using a bivariate spline approximation of order 3 of the original maps. Then, we cut from these maps patches of size in pixels $512\times 512$, with a step size of 256 pixels, such that two neighboring patches will have 50% of their pixels in common. In the following, these patches will be themselves tiled into 4 sub patches of size $256\times 256$ to perform a variability estimation. The choice of 6 ″ for the finest sampling allows us to exploit the 18 ″ resolution of these maps with mitigated sampling artifacts. The 9 ″, 12 ″, and 18 ″pixel sizes respectively correspond to a relative shrinking of the 6 ″ pixel maps with ratios of 1.5, 2, and 3, which allows to accommodate for the significant and quite uncertain range of distances of the different MCs in the HGBS. This entire procedure, which is illustrated in Fig. 5, and summarized in Table 1, leads to a total of $551$ patches with size $512\times 512$ pixels, each of which is then subdivided into four $256\times 256$ sub-patches.
@@
Figure 5: Illustration of the pre-processing of observational data. Details are given in Sec. 2.2.
Region Dist. Coord. # patches with resolution References
$d$ (pc) $l$ (°) $b$ (°) Tot. 6” 9” 12” 18”
Aquila 436 28 4 57 36 16 4 1 Ortiz-León et al. (2018)
Chamaeleon I 210 296 -16 28 20 6 2 0 Zucker et al. (2020)
Chamaeleon II III 160 302 -16 43 32 8 3 0 Zucker et al. (2020)
Corona Australis 160 0 -19 44 34 7 3 0 Zucker et al. (2020)
IC 5146 750 94 -5 4 4 0 0 0 Zucker et al. (2020)
Lupus I 182 339 17 17 12 4 1 0 Galli et al. (2013)
Musca 200 301 -9 6 6 0 0 0 Knude & Hog (1998)
Ophiuchus 139 353 17 61 42 12 6 1 Mamajek (2008)
Orion B 450 205 -14 88 61 19 7 1 Schlafly et al. (2014); Zucker et al. (2020)
Perseus 285 159 -20 44 32 10 2 0 Zucker et al. (2020)
Pipe Nebula 180 359 6 23 18 5 0 0 Zucker et al. (2020)
Polaris Flare 400 123 26 50 36 9 4 1 Schlafly et al. (2014); Zucker et al. (2020)
Serpens 436 31 4 46 33 9 4 0 Ortiz-León et al. (2018)
Taurus 145 170 -16 40 30 6 4 0 Yan et al. (2019)
Table 1: Summary of the properties of the different MCs studied in this paper, as well as their division into patches at different sampling. References in the rightmost column are for the distance estimates, which can be quite uncertain (uncertainties are given in the references quoted).
@@ 3.1 General methodology
We want to quantify the amount of information we can exploit from a given set of summary statistics on a given dataset. However, because we work with unlabeled data, we cannot use supervised frameworks, such as Fisher analysis, that have a label-based approach to quantify information. For instance, when working with simulations, each sampled data ${x_{i}}$ can be labeled by its corresponding physical parameters $\theta_{i}$. In our case, we have to deal with an ensemble of unlabeled maps, for which we have a priori no notion of distance between pairs.
@@
In this paper, we choose to rely on a notion of compatibility between patches, that can be estimated for any set of summary statistics. This approach allows to be quantitative, even in this unsupervised setting. Nevertheless, without supervision, it remains difficult to interpret a compatibility result between two patches for a given set of summary statistics: do the two patches actually have very similar properties, or do they have very distinct properties whose differences are not effectively caught by the statistics?
This difficulty can be partially overcome by using complementary sets of statistics. Indeed, if two patches are distinguished by a first set of statistics, this is sufficient to assess that they are in practice different, and therefore to highlight the degeneracies of another set of statistics. It is this comparative approach, which is all the more relevant when the panel of statistics compared is comprehensive, that we use in this paper.
@@
The informative power of a set of statistics strongly depends on the family of processes studied. For instance, we know that the empirical mean and power spectrum are sufficient statistics for stationary Gaussian fields (Cover, 1999), which is not the case for physical processes in general. In this paper, we will apply our approach to the dataset of observations defined above, as well as to numerical simulations and synthetic logFBM models. In addition, we note that this diagnostic can also be used to compare maps from two different datasets.
@@
In the rest of this section, we introduce the compatibility criterion, and explain how we extend it from a pair level to a dataset level. The set of statistics used in this paper will be presented in the following section.
Figure 6: Illustration of the proposed test to confront two sets of summary statistics, $\phi_{A}$ vs $\phi_{B}$, on their degeneracy level for a given dataset. Each star represents a pair of patches. Panel a: the presence of stars in the bottom right region, i.e., where $d_{\phi_{A}}\gg 1$ and $d_{\phi_{B}}\lessapprox 1$, reveals that some pairs of this dataset are identified by $\phi_{A}$ as incompatible but not by $\phi_{B}$: such pairs thus evidence degeneracies of $\phi_{B}$ lifted by $\phi_{A}$. Conversely, the presence of stars in the top left region evidences degeneracies of $\phi_{A}$ lifted by $\phi_{B}$. Hence, if all the points land mainly in the sub-diagonal part (panel b), this evidences that $\phi_{A}$ is better suited than $\phi_{B}$ to compare the pairs of this dataset. If on the contrary the points are spread both in the upper left and bottom right regions of the plot (panel c), this shows that both $\phi_{A}$ and $\phi_{B}$ are individually not sufficient to describe the processes of this dataset.
@@ 3.2 Statistical compatibility for a pair of patches
We want to measure a notion of compatibility, according to a given set of summary statistics $\phi$, between the two processes that generated the patches $(x_{i},x_{j})$. To estimate this $\phi$-compatibility, we need to make a number of simplifications, given the low-data regime, which will bring us back to a simplified case of statistical hypothesis testing. To do so, we tile each patch $x_{i}$ into 4 sub patches $\{x_{i}^{(l)}\}_{1\leq l\leq 4}$, as illustrated in Fig. 5. We then compute the statistics at this sub patch level: $\{\phi(x_{i}^{(l)})\}_{1\leq l\leq 4}$ and we assume that these random variables can be considered as independent samples of the same distribution, that we furthermore model as multivariate normal distribution [^5] This assumption is not far from being true when the law of large numbers can be applied to the distribution of $\phi(x)$. This is for instance the case when $\phi(x)$ is defined as an average over the image pixels of a certain local distortion (such as filtering, possibly non linear) $\phi_{loc}(x)$ of $x$: $\phi(x)=\langle\phi_{loc}(x)[\@vec{u}]\rangle_{\@vec{u}}$ and when the process $\phi_{loc}(x)$ is stationary with a correlation length that is small compared to the image length. of mean $\mu_{i}$ and variance $\Sigma_{i}$:
@@
Under these assumptions, the problem boils down to testing the compatibility between the two normal distributions, i.e., to test the hypothesis: $\mu_{i}=\mu_{j}$ and $\Sigma_{i}=\Sigma_{j}$. In our case, however, we have to estimate this compatibility from very few samples, most of the time fewer than the dimension of the vector of statistics $\phi$. We thus choose to focus only on testing if the means $\mu_{i}$ and $\mu_{j}$ are statistically compatible, but not $\Sigma_{i}$ and $\Sigma_{j}$, a problem known as the multivariate two-sample mean test. The most widespread test statistic for this problem is the Hotelling’s two-sample $T^{2}$-statistic, a multivariate extension of Student’s $t$-test (Hotelling, 1931). However, this test statistic requires to invert an estimation $S$ of the full covariance matrix $\Sigma_{i}+\Sigma_{j}$, which is usually intractable in our low data regime.
To overcome this, Srivastava & Du (2008) proposed a test statistic based only on the diagonal $D_{S}$ of the covariance estimator $S$, and on the trace of the square of its associated correlation matrix $R\equiv D_{S}^{-1/2}SD_{S}^{-1/2}$:
@@
where $\hat{\mu}_{i}$ is the estimator of $\mu_{i}$ obtained through an average over the $4$ sub-patches $x_{i}^{(l)}$, $\alpha$ is an overall factor to normalize the variance of $d^{2}_{\phi}$, and $\beta$ is a debiasing term. We emphasize that, through a dependence on $tr(R^{2})$, the $\alpha$ factor accounts, at least partially, for the correlation structure of $\phi$. More details about these terms are given in appendix C.
Under the assumption that $\Sigma_{i}=\Sigma_{j}$, the $d_{\phi}^{2}(x_{i},x_{j})$ test statistic has a variance of order unity. Thus, when it is much larger than 1, $\mu_{i}$ $\mu_{j}$ cannot be considered compatible:
$d_{\phi}^{2}(x_{i},x_{j})\gg 1\implies\mu_{i}\text{ \& }\mu_{j}\text{ incompatible},$ (3)
whereas when it is of the order of 1 or less, it is not possible to detect a discrepancy between the two means $\mu_{i}$ and $\mu_{j}$ with the available amount of data:
$d_{\phi}^{2}(x_{i},x_{j})\lessapprox 1\implies\mu_{i}\text{ \& }\mu_{j}\text{ not incompatible based on}$ (4)
@@ 3.3 Comparing summary statistics on a dataset
By extending the $\phi$-compatibility test introduced above, we set up a comparison between two sets of summary statistics $\phi_{A}$ and $\phi_{B}$ on a given dataset. This comparison consists in studying whether each set of summary statistics has degeneracies that the other set can lift. These degeneracies are evidenced by the presence in the dataset of pairs of maps that are clearly incompatible for $\phi_{A}$, but for which $\phi_{B}$ detects no incompatibility, or vice versa. To be relevant, this comparison requires that the dataset contain maps that actually have different properties, which will be the case for the data sets studied later.
@@
  • Step 1: for every patch $x_{i}$ of the dataset, compute the statistics $\phi_{A}$ and $\phi_{B}$ of its corresponding sub-patches $\{x_{i}^{(l)}\}_{l}$.
@@
The resulting scatter plot can then be used to detect whether there are some pairs of processes in the dataset that are degenerate for $\phi_{A}$ but lifted by $\phi_{B}$ (or vice versa), as explained in Fig. 6. In such plots, the presence of points in the bottom right region, i.e., where $d_{\phi_{A}}\gg 1$ and $d_{\phi_{B}}\lessapprox 1$, reveals that some pairs of this dataset are identified by $\phi_{A}$ as incompatible but not by $\phi_{B}$: such pairs thus evidence degeneracies of $\phi_{B}$ lifted by $\phi_{A}$ (panel a). Conversely, the presence of points in the top left region evidences degeneracies of $\phi_{A}$ lifted by $\phi_{B}$. Hence, if all the points land mainly in the sub-diagonal part (panel b), this evidences that $\phi_{A}$ is better suited than $\phi_{B}$ to compare the pairs of this dataset. If on the contrary the points are spread both in the upper left and bottom right regions of the plot (panel c), this shows that both $\phi_{A}$ and $\phi_{B}$ are individually not sufficient to describe the processes of this dataset.
@@ 4 Summary statistics
Note that for some usual statistics, we consider their logarithms instead of their raw values. For instance we use $\log\,\langle x\rangle_{\@vec{u}}$ instead of $\langle x\rangle_{\@vec{u}}$. This is possible when we work with positive-valued statistics. We made this choice to better fit the Gaussianity assumption given Eq. (1). In addition, we found that logarithms expressed more discriminatory power with our diagnostic. See appendix D for more details. If no precision is made, $\log$ designates the logarithm in base 10.

### v2

@@ 2.2 Subsampling and tiling
Figure 4: Illustration of the preprocessing of observational data. Details are given in Sect. 2.2.
The presence of physical processes such as gravity in the dynamics of MCs implies spatial variations of their statistical properties, which prevents us from modeling them as stationary stochastic processes. For instance, they often exhibit strong local overdensities, while being surrounded by diffuse borders. This inhomogeneity makes it difficult to compare such objects as a whole, as well as to properly estimate their statistical properties, when seen as realizations of a random process, as these properties will, for instance, depend heavily on the identification of the cloud boundary, or on the precise definition of what a cloud is. Furthermore, variance estimates for random processes generally rely on a homogeneity assumption to estimate the intrinsic variability of a process from its spatial variations. More broadly, the estimation of the statistical moments and properties of a process is usually based on spatial averaging, which assumes the homogeneity of the sample.
To avoid this problem, we choose to restrict the comparison to local patches of these MCs, assuming statistical homogeneity within each individual patch. This requires the identification of a characteristic stationarity length over which it can be assumed that the statistical properties of the cloud do not vary significantly. Different patches extracted from the same cloud may have different statistical properties, which will be representative of the inhomogeneity of the cloud as a whole. This encourages us to use small patches. However, an additional difficulty is that these patches must be large enough to make statistical estimates. In particular, variance estimates require sampling beyond the correlation length of the process under study.
In this paper, we have chosen to cut patches with angular sizes ranging from $0.85$° to $2.55$°, which correspond respectively to 333\,3pc and 999\,9pc for a cloud at a typical distance of 200200200\,200pc. We believe this represents a good compromise with regard to the trade-off mentioned above. However, we are aware that there may not be an ideal solution, especially as we have no precise estimate of either the stationarity length or the correlation length. This means that we cannot rule out potential confusion between local intrinsic variability and large-scale variations in the statistical properties of the clouds. However, we believe that, for lack of a better solution, this does not undermine the relevance of this study.
More precisely, we make multiple versions of the original (18″ resolution, 3″ pixel size) column density maps at different pixel sizes (6″, 9″, 12″, and 18″). The subsampling is done using a bivariate spline approximation of order three of the original maps. Then, we cut from these maps patches of size in pixels $512\times 512$, with a step size of 256 pixels, such that two neighboring patches have 50% of their pixels in common. In the following, these patches are tiled into four subpatches of size $256\times 256$ to perform a variability estimation. The choice of 6″ for the finest sampling allows us to exploit the 18″ resolution of these maps with mitigated sampling artifacts. We note furthermore that the smallest scales that we probe with the two-point-based statistics presented in Sect. 4.2 remain above 24″. This prevents the finite 18″ resolution to significantly affect the results. The 9″, 12″, and 18″ pixel sizes respectively correspond to a relative shrinking of the 6″ pixel maps with ratios of 1.5, 2, and 3, which allows us to accommodate for the significant and quite uncertain range of distances of the different MCs in the HGBS. This entire procedure, which is illustrated in Fig. 4, and summarized in Table 1, leads to a total of $551$ patches with size $512\times 512$ pixels, each of which is then subdivided into four $256\times 256$ subpatches.
@@ 3.1 General methodology
We want to quantify and compare the informative power of different sets of statistics on a given dataset, that is, their ability to distinguish between patches of different physical properties. However, because we work with unlabeled data, we cannot use supervised frameworks, such as Fisher analysis, that have a label-based approach to quantify information. For instance, when working with simulations, each data sample ${x_{i}}$ can be labeled by its corresponding physical parameters $\theta_{i}$. In our case, we have to deal with an ensemble of unlabeled maps, for which we have a priori no notion of similarity between pairs.
@@
In this paper, we choose to rely on a notion of compatibility between patches, computed from their summary statistics, and that can be estimated for any set of statistics. This approach allows us to be quantitative, even in such an unsupervised setting. Nevertheless, without supervision, it remains difficult to interpret a compatibility result between two patches for a given set of summary statistics. Indeed, it is unclear whether the two patches actually have similar properties or if they possess genuinely distinct properties whose differences are not effectively captured by the statistics.
To overcome this difficulty, we compare the results obtained from complementary sets of statistics. Indeed, we assume that if two patches are distinguished by a first set of statistics, this is sufficient to assess that these patches have different physical properties666 Indeed, this implication, from statistical to physical incompatibility, is ensured for patches that undergo stationary dynamics, both in space and time. However, for dense molecular clouds, this assumption is not entirely satisfied when nonstationary processes such as self-gravity come into play. A potential caveat could then be to find incompatibilities between summary statistics which are not particularly relevant from a physical point of view. While we do not further discuss this issue in the present paper, we think it could deserve further work. and therefore to highlight the degeneracies of another set of statistics. It is this comparative approach, which is all the more relevant when the panel of statistics compared is comprehensive, that we use in this paper.
@@
We also emphasize that the informative power of a set of statistics strongly depends on the family of processes studied. For instance, we know that the empirical mean and PS are sufficient statistics for stationary Gaussian fields (Cover & Thomas, 2006), which is not the case for physical processes in general. In this paper, we apply our approach to the dataset of observations defined above as well as to numerical simulations and synthetic logFBM models. In addition, we note that this diagnostic can also be used to compare maps from two different datasets.
@@
In the rest of this section, we introduce the compatibility criterion, and explain how we extend it from a level of pair of patches to a dataset level. The sets of statistics used in this paper are presented in the following section.
@@ 3.2 Statistical compatibility for a pair of patches
We aim to measure a notion of compatibility, according to a given set of summary statistics $\phi$, between the two processes that generated the patches $(x_{i},x_{j})$. To estimate this $\phi$-compatibility, we need to make a number of simplifications, given the low data regime, that will bring us back to a simplified case of statistical hypothesis testing. To do so, we tiled each patch $x_{i}$ into four subpatches $\{x_{i}^{(l)}\}_{1\leq l\leq 4}$, as illustrated in Fig. 4. We then computed the statistics at this subpatch level, $\{\phi(x_{i}^{(l)})\}_{1\leq l\leq 4}$, and we assumed that these random variables can be considered as independent samples of the same distribution, which we furthermore modeled as a multivariate normal distribution [^7] This assumption is not far from being true when the law of large numbers can be applied to the distribution of $\phi(x)$. This is for instance the case when $\phi(x)$ is defined as an average over the image pixels of a certain local distortion (such as filtering, possibly nonlinear) $\phi_{loc}(x)$ of $x$: $\phi(x)=\langle\phi_{loc}(x)[\@vec{u}]\rangle_{\@vec{u}}$ and when the process $\phi_{loc}(x)$ is stationary with a correlation length that is small compared to the image length. of mean $\mu_{i}$ and variance $\Sigma_{i}$:
@@
Under these assumptions, the problem of estimating the compatibility between the distributions of $\phi(x_{i}^{(l)})$ and $\phi(x_{j}^{(l)})$ boils down to testing the compatibility between the two normal distributions, that is, to test the hypothesis: $\mu_{i}=\mu_{j}$ and $\Sigma_{i}=\Sigma_{j}$. In our case, however, we have to estimate this compatibility from very few samples, most of the time fewer than the dimension of the vector of statistics $\phi$. Thus, we choose to focus only on testing if the means $\mu_{i}$ and $\mu_{j}$ are statistically compatible, but not $\Sigma_{i}$ and $\Sigma_{j}$, a problem known as the multivariate two-sample mean test. The most widely used test statistic for this problem is Hotelling’s two-sample $T^{2}$-statistic, a multivariate extension of Student’s $t$-test (Hotelling, 1931). However, this test statistic requires to invert an estimation $S$ of the full covariance matrix $\Sigma_{i}+\Sigma_{j}$, which is usually intractable in our low data regime.
To overcome this, Srivastava & Du (2008) proposed a test statistic based only on the diagonal $D_{S}$ of the covariance estimator $S$ (defined in Appendix B), and on the trace of the square of its associated correlation matrix $R\equiv D_{S}^{-1/2}SD_{S}^{-1/2}$:
@@
where $\hat{\mu}_{i}$ is the estimator of $\mu_{i}$ obtained through an average over the four subpatches $x_{i}^{(l)}$, $\alpha$ is an overall factor to normalize the variance of $d^{2}_{\phi}$, and $\beta$ is a debiasing term. We emphasize that, through a dependence on $\operatorname{tr}(R^{2})$, the $\alpha$ factor accounts, at least partially, for the correlation structure of $\phi$. Further details about these terms are provided in Appendix B.
Under the assumption that $\Sigma_{i}=\Sigma_{j}$, the $d_{\phi}^{2}(x_{i},x_{j})$ test statistic has a variance of order unity. Thus, when it is much larger than one, $\mu_{i}$ and $\mu_{j}$ cannot be considered compatible:
$d_{\phi}^{2}(x_{i},x_{j})\gg 1\implies\mu_{i}\text{ and }\mu_{j}\text{ are incompatible};$ (3)
whereas when it is of the order of one or less, it is not possible to detect a discrepancy between the two means $\mu_{i}$ and $\mu_{j}$ with the available amount of data:
$d_{\phi}^{2}(x_{i},x_{j})\lessapprox 1\implies\mu_{i}\text{ and }\mu_{j}\text{ are not incompatible based on}$ (4)
@@ 3.3 Comparing summary statistics on a dataset
Figure 5: Illustration of the proposed test to confront two sets of summary statistics, $\phi_{A}$ versus $\phi_{B}$, on their degeneracy level for a given dataset. Each star represents a pair of patches. Panel a: the presence of stars in the bottom right region, i.e., where $d_{\phi_{A}}\gg 1$ and $d_{\phi_{B}}\lessapprox 1$, reveals that some pairs of this dataset are identified by $\phi_{A}$ as incompatible but not by $\phi_{B}$: such pairs thus evidence degeneracies of $\phi_{B}$ lifted by $\phi_{A}$. Conversely, the presence of stars in the top left region evidences degeneracies of $\phi_{A}$ lifted by $\phi_{B}$. Hence, if all the points land mainly in the sub-diagonal part (panel b), this evidences that $\phi_{A}$ is better suited than $\phi_{B}$ to compare the pairs of this dataset. If on the contrary the points are spread both in the upper left and bottom right regions of the plot (panel c), this shows that both $\phi_{A}$ and $\phi_{B}$ are individually not sufficient to describe the processes of this dataset.
By extending the $\phi$-compatibility test introduced above, we set up a comparison between two sets of summary statistics $\phi_{A}$ and $\phi_{B}$ on a given dataset. This comparison consists in studying whether each set of summary statistics has degeneracies that the other set can lift. These degeneracies are evidenced by the presence in the dataset of pairs of maps that are clearly incompatible for $\phi_{A}$, but for which $\phi_{B}$ detects no incompatibility, or vice versa. For this comparison to be relevant, the dataset must contain maps with genuinely different properties, which is the case for the sets studied below.
@@
  • Step 1: For every patch $x_{i}$ of the dataset, compute the statistics $\phi_{A}$ and $\phi_{B}$ of its corresponding subpatches $\{x_{i}^{(l)}\}_{l}$.
@@
The resulting scatter plot can then be used to detect whether there are some pairs of processes in the dataset that are degenerate for $\phi_{A}$ but not for $\phi_{B}$ (or vice versa), as explained in Fig. 5. In such plots, the presence of points in the bottom right region, that is, where $d_{\phi_{A}}\gg 1$ and $d_{\phi_{B}}\lessapprox 1$, reveals that some pairs of this dataset are identified by $\phi_{A}$ as incompatible but not by $\phi_{B}$: such pairs thus evidence degeneracies of $\phi_{B}$ lifted by $\phi_{A}$ (panel a). Conversely, the presence of points in the top left region evidences degeneracies of $\phi_{A}$ lifted by $\phi_{B}$. Hence, if all the points land mainly in the subdiagonal part (panel b), this evidences that $\phi_{A}$ is better suited than $\phi_{B}$ to compare the pairs of this dataset. If, on the contrary, the points are spread both in the upper left and bottom right regions of the plot (panel c), this shows that both $\phi_{A}$ and $\phi_{B}$ are individually insufficient to describe the processes of this dataset.
@@ 4 Summary statistics
We note that the authors consider the logarithms of some typical statistics, instead of their raw values. For instance, we use $\log\,\langle x\rangle_{\@vec{u}}$ instead of $\langle x\rangle_{\@vec{u}}$. This is possible when we work with positive-valued statistics. We made this choice to better fit the Gaussianity assumption given in Eq. (1). In addition, $\phi(x)$ and $\log\,\phi(x)$ should hold the same amount of information as one can be retrieved from the other, see Appendix C for more details. For readability, we name in the following ”$\phi$” instead of ”log of $\phi$”. We note, however, that neither $\phi(x)$ nor $\log(\phi(x))$ should be mistaken for $\phi(\log x)$, and we name the latter ”$\phi$ of log”. If no precision is made, $\log$ designates the logarithm in base ten.